OpenStudy (anonymous):

Avery has a total of $12.50 in her piggy bank at home. Each week she takes out $1.25 to ride the city bus. Write an equation representing the balance, C, in the piggy bank, where w is the number of weeks.

ganeshie8 (ganeshie8):

Since the starting amount is \($12.50\), the amount in her piggy bank in \(\text{week 0}\) is \($12.50\)

ganeshie8 (ganeshie8):

Can you guess the amount in her piggy bank during \(\text{week 1}\) ?

OpenStudy (anonymous):

11.25

ganeshie8 (ganeshie8):

Yes! because she takes out \(1.25\) : \[\large 12.50-\color{red}{1.25} = 11.25\]

ganeshie8 (ganeshie8):

can you guess the amount in her piggy band during \(\text{week 2}\) ?

OpenStudy (anonymous):

10

ganeshie8 (ganeshie8):

Correct! you must have worked it by subtracting \(1.25\) again : \[\large 12.50-\color{red}{1.25} -\color{red}{1.25} = 10\]

OpenStudy (anonymous):

i know how to subtract. i just need help with how i write the equation. I know that it begins like "C=......"

ganeshie8 (ganeshie8):

Notice the pattern ? can you find the amount in her piggy bank during \(\text{week} 8\) ?

ganeshie8 (ganeshie8):

you just need to subtract \(\color{red}{1.25}\) for \(8\) times right

ganeshie8 (ganeshie8):

so the amount in the bank after \(w\) weeks is given by : \[C ~~=~~12.50~-~\color{red}{1.25}w\]

OpenStudy (anonymous):

Those are my choices C = 1.25w – 12.50 C = 1.25w + 12.50 C = –1.25w + 12.50 C = –1.25w – 12.50

OpenStudy (anonymous):

Would i just move it around????

OpenStudy (anonymous):

I thought that it was C=12.50-1.25w but thats not up there.

ganeshie8 (ganeshie8):

Yes \[C ~~=~~12.50~-~\color{red}{1.25}w\] is same as \[C ~~=~-~\color{red}{1.25}w~+~12.50\]

OpenStudy (anonymous):

ohhhhhhh okay. haha. thats all i really needed to know. Thank you.

ganeshie8 (ganeshie8):

yw!